1.              Assume you are managing a food processing plant in Ethiopia. The demand function for one of your product is given as Qd=50-2p. (8 marks)

a)       Find the point price elasticity if price is 15 ETB? Is it elastic or inelastic?

b)      How do you interpret the elasticity result?

c)       In order to get more revenue what will be your recommendation. Is it to increase price or decrease price? Why?

d)      Describe at least four determinants of the price elasticity demand for the food product?


1
Expert's answer
2021-06-15T12:43:18-0400

a) First we will find Qd when price equal to 15

"Qd = 50 - 2p"

"Qd = 50 - 2\u00d715 = 50 - 30 = 20"

Point price elasticity  "= \frac{∆Qd}{\Delta P}\times \frac{P}{Q_d}"

Now we will find "\frac{∆Q_d}{\Delta p}"

"\frac{∆Q_d}{\Delta p}=-2"

Point price elasticity "= -2 × \frac{p}{Qd} = -2 × \frac{15}{20}"

"= -2 \u00d7 0.75 = - 1.5"

In absolute terms point price elasticity is 1.5. The point price elasticity in absolute terms is greater than 1 which means the demand is elastic.

b) -1.5 elasticity means that when price increases by 1 ETB, the quantity demanded will reduce by 1.5 units. Elasticity measures the responsiveness in quantity demanded due to change in price of a good.

c) Revenue equals price × quantity. In order to increase revenue, price cannot be increased as in this case the demand is elastic and due to which quantity demanded will fall by more than the increase in price. While a small decrease in price might help to get more revenue as the quantity demanded will increase by more than the price decrease.

d)Substitutes for the product: better substitutes, the more elastic.

Proportion of expenditures on specified good relative to income: the larger the expenditure on the specific good relative to ones budget, the more elastic.

Product is a luxury or a necessity: less necessary the item the more elastic.

Longer the time period involved in making the purchase, the more elastic.